OpenStudy (anonymous):

Which is the correct classification of the system of equations below? 12x-3y=9 y-4x=-3 A)parallel B)coincident C)intersecting

OpenStudy (zzr0ck3r):

can you solve the system?

OpenStudy (zzr0ck3r):

you either get 1, none, or infinite solutions

OpenStudy (zzr0ck3r):

1 implies intersecting, infinite implies they are the same, 0 implies they have the same slope but never touch

OpenStudy (anonymous):

So is it coincident?

OpenStudy (anonymous):

the second equation is \(-4x +y = -3 \Longleftrightarrow -12x + 3y = -12 \Longleftrightarrow 12x - 3y = 12 \). these equations can have solutions in a universe where \(9 = 12.\)